A woman is now 30 years older than her son. 15 years ago, she was twice as old. What are the present ages of the woman and her son?
Answer by Maths68(1474)

Let
Present age of woman = w years old
Present age of son = s years old
woman is now 30 years older than her son
w=s+30.........(1)
15 years ago
woman was = w-15 years old
son was = s-15 years old
Then
she was twice as old.
w-15=2(s-15)
w-15=2s-30
w=2s-30+15
w=2s-15...........(2)
Put the value of w from (1) to (2)
s+30=2s-15
s-2s=-15-30
-s=-45
s=45
Put the value of s in (1)
w=s+30.........(1)
w=45+30
w=75

Present age of woman = w = 75 years old
Present age of son = s = 45 years old